OpenStudy (anonymous):

Decide what number must be added to the binomial to make a perfect square trinomial: x^(2) - 6x + ___

OpenStudy (anonymous):

9

OpenStudy (anonymous):

(x-3)^2 ax^2+2ab+b^2 a=1, so 6/2=3 so b=3 b^2=9